OpenStudy (anonymous):

Explain why 3^6 x 5^4 is a perfect square

OpenStudy (amistre64):

becasue when you take the sqrt of it, nothing is left under the radical

OpenStudy (amistre64):

the exponents are even so they are divisible by 2

OpenStudy (amistre64):

sqrt(3^6 5^4) = 3^(6/2) 5^(4/2) = 3^3 5^2

OpenStudy (sasogeek):

it is a perfect square because the square root of the result is an integer

OpenStudy (anonymous):

amistre64 haha.. right :)

OpenStudy (anonymous):

yes, the square root is an integer.

OpenStudy (amistre64):

perfect squares need not be integers tho ... just unradicalized :) if i recall it right

OpenStudy (sasogeek):

numbers without decimal places...

OpenStudy (amistre64):

http://www.mathwords.com/p/perfect_square.htm